# Counterfactual Fairness by Combining Factual and Counterfactual Predictions

### v1

@@ Abstract
In high-stake domains such as healthcare and hiring, the role of machine learning (ML) in decision-making raises significant fairness concerns. This work focuses on Counterfactual Fairness (CF), which posits that an ML model’s outcome on any individual should remain unchanged if they had belonged to a different demographic group. Previous works have proposed methods that guarantee CF. Notwithstanding, their effects on the model’s predictive performance remains largely unclear. To fill in this gap, we provide a theoretical study on the inherent trade-off between CF and predictive performance in a model-agnostic manner. We first propose a simple but effective method to cast an optimal but potentially unfair predictor into a fair one without losing the optimality. By analyzing its excess risk in order to achieve CF, we quantify this inherent trade-off. Further analysis on our method’s performance with access to only incomplete causal knowledge is also conducted. Built upon it, we propose a performant algorithm that can be applied in such scenarios. Experiments on both synthetic and semi-synthetic datasets demonstrate the validity of our analysis and methods.
@@ 1 Introduction
Machine learning (ML) has been widely used in high-stake domains such as healthcare (Daneshjou et al., 2021), hiring (Hoffman et al., 2018), criminal justice (Brennan et al., 2009), and loan assessment (Khandani et al., 2010), bringing with it critical ethical and social considerations. A prominent example is the bias observed in the COMPAS tool against African Americans in recidivism predictions (Brackey, 2019). This issue is particularly alarming in an era where large-scale deep learning models, commonly trained on noisy data from the internet, are increasingly prevalent. Such models, due to their extensive reach and impact, amplify the potential for widespread and systemic biases. This increasing awareness underscores the need for ML practitioners to integrate fairness considerations into their work, extending their focus beyond merely maximizing prediction accuracy (Bolukbasi et al., 2016; Calders and Verwer, 2010; Dwork et al., 2012; Grgic-Hlaca et al., 2016; Hardt et al., 2016). Various fairness notions have been developed, ranging from group-level measures such as group parity (Hardt et al., 2016) to individual-level metrics (Dwork et al., 2012). Recently, there has been a growing interest in approaches based on causal inference, particularly in understanding the causal effects of sensitive attributes such as gender and age on decision-making (Chiappa, 2019; Galhotra et al., 2022; Khademi et al., 2019). This has led to the proposal of Counterfactual Fairness (CF), which states that prediction for an individual in hypothetical scenarios where their sensitive attributes differ should remain unchanged (Kusner et al., 2017). As an individual-level notion agnostic to the choice of similarity measure (Kusner et al., 2017; Rosenblatt and Witter, 2023a), CF has recently gained traction (Anthis and Veitch, 2024; Nilforoshan et al., 2022; Makhlouf et al., 2022; Rosenblatt and Witter, 2023b).
@@
To take a first step in understanding this problem, we conduct a formal quantitative study on the inherent trade-off between Counterfactual Fairness (CF) and predictive performance, which quantifies the predictive power that has to be given up for satisfying the fairness (i.e., perfect CF) criteria(Zhao and Gordon, 2022; Xian et al., 2023). Towards this goal, we first prove that the optimal predictor under the fairness constraint can be achieved by combining factual and counterfactual predictions using a (potentially unfair) optimal predictor. To the best of our knowledge, this is the first counterfactually fair predictor that is proven optimal. Next, we quantify the excess risk between the optimal predictor with and without CF constraints. This quantity sheds lights on the best possible model, in terms of predictive performance, under the stringent notion of perfect CF. To take scenarios with incomplete causal knowledge (e.g. unknown causal graph or model) in consideration, we further study the CF and predictive performance degradation caused by imperfect counterfactual estimations. Inspired by our theoretical findings, we first propose a plugin method that leverages a (potentially unfair) pretrained model to achieve optimal fair prediction and further propose a practical way to achieve good empirical performance under counterfactual estimation inaccuracies.
We summarize our contributions as follows:
@@
    To the best of our knowledge, we propose the first CF method that is provably optimal in terms of predictive performance under perfect CF and characterize their inherent trade-off, which applies to all CF methods.
@@
  3. 3.
    We empirically demonstrate that our proposed CF methods outperform existing methods in both the full and incomplete causal knowledge settings.
@@ 2 Preliminaries
2.1 Notation
We use capital letter to represent random variables and lowercase letters to represent the realization of random variables. Now we define a few variables that will be considered in this work. $A$ represents the sensitive attribute of an individual (e.g. gender), $Y$ represents the target variable to predict, $X$ represents observed features other than $A$ and $Y$, and $U$ represents unobserved confounding variables which are not caused by any observed variables while $a,y,x,u$ represent their realization respectively.
2.2 Counterfactual
In this work, we use the framework of Structrual Causal Models (SCMs) (Pearl, 2009). A SCM is a triplet $\mathcal{M}=(\bf{U},\bf{V},\mathcal{F})$ where $\bf{U}$ represent exogenous variables (factors outside the model), $\bf{V}$ represent endogenous variables, and $\mathcal{F}$ contains a set of functions $F_{i}$ that maps from $U_{i}$ and $Parent(V_{i})$ to $V_{i}$.
A counterfactual query asks a question like: what would the value of $Y$ be if $A$ had taken a different value given certain observations $x$. For example, given that a person is a woman and given everything we observe about her performance in an interview, what is the probability of her getting the job if she had been a man? More formally, given a SCM, a counterfactual query can be written as $P(Y_{A=a}|W=w)$. Here $W=w$ is the evidence and $A=a$ in the subscript represents the intervention on $A$. For the general procedure to estimate counterfactual, please refer to Pearl (2009).
2.3 Counterfactual Fairness
Built upon the framework above, we focus on Counterfactaul Fairness (CF), which requires the predictors to be fair among factual and counterfactual samples. More formally, it is defined as below
@@ 2.3 Counterfactual Fairness
$A$$U$$Y$X
Figure 1: Causal graph. $A$ represents sentitive attribute, $Y$ represents the target variable, $U$ represents latent confounders, $X$ represents observed features other than $A$ and $Y$.
@@ 3.1 Problem Setup
We assume that all data we have is generated by a causal model (Pearl, 2009) and we consider the representative causal graph shown in Figure 1 that has been widely adopted in the literature (Grari et al., 2023; Kusner et al., 2017; Zuo et al., 2023) For simplicity, we omit the exogenous noise variables corresponding to $A$ and $Y$ in the graph. Our analysis is presented based on binary $A\in\{0,1\}$ given its pivotal importance in the literature (Pessach and Shmueli, 2023) and for the sake of presentation clearness, but our analysis and our method can be naturally extended on multi-class $A$. We first state the main assumptions needed in this section.
@@
While the invertibility assumption might be restrictive in certain scenarios, it simplifies the theoretical analysis and has been adopted in recent works on counterfactual estimation (Nasr-Esfahany et al., 2023; Zhou et al., 2024). Further, we empirically validate the effectiveness of our method after relaxing this assumption in the experiment section. To facilitate our discussion, we first define $F_{X}$ as the mapping between $X$ and $(U,A)$, i.e., $X=F_{X}(U,A)$. According to our second assumption, $F_{X}(\cdot,a)$ is an invertible function, i.e., $\exists{F_{X}^{*}}^{-1},{F_{X}^{*}}^{-1}(x,a)={F_{X}^{*}}^{-1}(F_{X}^{*}(u,a),a)=u,\forall(x,a)$. This assumption simplifies the counterfactual estimation of $X$ for different values of $A$ into a deterministic function. In our context, counterfactual estimation specifically refers to the estimation of $X$. We introduce the concept of the counterfactual generating mechanism (CGM), denoted as $x_{a\to a^{\prime}}=G(x,a,a^{\prime})$. All proof can be found in Appendix A.
Given the setup, the following lemma characterizes the perfect CF constraint on $\phi$.
@@
The proof is straightforward from the definition of TE. Notably, this lemma helps disambiguate the question whether counterfactual fairness is a distribution- or individual-level requirements as raised in Plecko and Bareinboim (2022). In our setup, they are equivalent due to invertibility between $X$ and $U$ given $A$.
@@ 3.2 Optimal Counterfactual Fairness and Inherent Trade-off
Given the complete knowledge of the causal model, it is viable to satisfy the perfect CF constraint (Kusner et al., 2017; Zuo et al., 2023). These methods are known to result in empirical degradation of ML models performance, raising critical concerns about the fairness-utility trade-offs. However, it is still unknown to what extent the ML model performance has to be affected in order to achieve perfect CF. In this section we provide a formal study on this to close the gap.
Our solution consists of two steps. First, we propose a simple yet effective method that is provably optimal under the constraint of perfect CF. Next, we characterize the inherent trade-off between CF and predictive performance by checking the excess risk of this Bayes optimal predictor. Our result shows that the inherent trade-off is dominated by the dependency between $Y$ and $A$, echoing with previous analysis on non-causal based fairness notions (Chzhen et al., 2020; Xian et al., 2023). For brevity, we refer to a Bayes optimal predictor as “optimal”, and a model that satisfies perfect CF as “fair”.
@@
This result suggests that, if we have to access to ground truth counterfactuals, a simple algorithm using an (potentially) unfair model could achieve strong fairness and accuracy. Remarkably, to our best knowledge, this is the first predictor that proves to achieve the optimality while being fair. Built upon the above result, we are ready to characterize the inherent trade-off between CF and model performance by the following theorem.
@@
Remarkably, the excess risks are completely characterized by the inherent dependency between $Y$ and $A$ as determined by the underlying causal mechanism, similar to non-causal based group fairness (Chzhen et al., 2020; Xian et al., 2023). Moreover, they lower bound the excess risk of all possible predictors one may apply in order to achieve perfect CF.
@@ 3.3 Method with Incomplete Causal Knowledge
In this section, we aim at addressing CF in the scenario where causal knowledge is limited. Inspired by Theorem 3.3, we first present a simple plugin method as summarized in Algorithm 1.
@@
For regression tasks, $\hat{\mu}$ is the final output, and for classification tasks, $\hat{\mu}$ is equivalent to the probability of $Y=1$, i.e., $p(Y=1|X=x,A=a)=\mathbb{E}[Y|X=x,A=a]$. It is note worthy that PCF is agnostic to the training of predictor $\phi$ that can be determined by the user freely. In fact, with access to the oracle CGM $G^{*}$, then PCF would achieve perfect CF as proved in the next result.
@@
Note that the this proposition only requires access to ground truth $G^{*}$ and holds valid for any pretrained predictor $\phi$. If $\phi$ is further accurate, then the corresponding PCF is able to achieve high accuracy as well, which is empirically validated in the experiments.
@@ 3.3.1 Given estimated $G$
Acquiring counterfactuals in practice can be a challenging task and could lead to estimation errors. In this section we provide a theoretical analysis on this. Specifically, theorem below bounds the TE and excess risk due to the use of estimated counterfactuals.
@@
Given an optimal predictor $\phi^{*}(x,a)$. Suppose it is L-lipschitz continuous in $x$, and the counterfactual estimation error is bounded, i.e.,
@@
for some $\varepsilon\geq 0$, $G^{*}$ and $\hat{G}$ are the ground truth and estimated CGMs respectively. Then, the total effect (TE) of Algorithm 1 based on $\hat{G}$ is bounded by $L\varepsilon$. Moreover, for squared $L_{2}$ loss  [^2] We conjecture the error bound with cross-entropy loss follows a similar form and leave the more careful investigation as the future work., the excess risk is bounded by $\sigma_{A}^{2}L^{2}\varepsilon^{2}+2\sigma_{A}^{2}L\varepsilon\mathbb{E}_{U}[|\mathbb{E}[Y\mid U=u,A=1]-\mathbb{E}[Y\mid U=u,A=0]|]$.
Note that $\sigma_{A}$ and $\mathbb{E}_{U}[|\mathbb{E}[Y\mid U=u,A=1]-\mathbb{E}[Y\mid U=u,A=0]|]$ are inherent characteristic of the underlying mechanism and is independent of the counterfactual estimation. This suggests that if the counterfactuals are not too far away and $\phi^{*}$ is smooth, then fairness and prediction performance will not be affected significantly.
@@ 3.3.2 Given estimated $G$ and $\phi$
The proof is straightforward since the original proof in Theorem 3.6 does not use any characteristic of optimality. To achieve good predictive performance, a natural approach is to train $\phi$ on the observed data via Empirical Risk Minimization (ERM), which should fit the predictor well given sufficient samples and a reasonable predictor class. However, ERM can only approximate Bayes optimality within the support of the training data. Outside this support, its performance can deteriorate significantly, as extensively studied in areas such as Domain Adaptation (Farahani et al., 2021) and Domain Generalization (Zhou et al., 2022). Consequently, when integrated with an approximate $G$, we may encounter the issue where $p(\hat{X}_{A=1-a}|A=a)\neq p(X|A=1-a)$, inducing a distribution shift problem. To mitigate this, we suggest improving $\phi$ on the estimated counterfactual distribution. More formally we define the following objective called Counterfactual Risk Minimization (CRM)
@@
This can be achieved either by augmenting the original training dataset or by fine-tuning with counterfactual samples. The choice between training from scratch or fine-tuning depends on the scale of the experiment and computational constraints. One important note is that the $Y$ corresponding to the estimated counterfactual should remain the same as the $Y$ of the factual samples. The motivation is that, given the formulation in Theorem 3.3, we conjecture that training on a mixture of factual and counterfactual samples while keeping $Y$ the same is optimal. We leave a more thorough investigation of this for future work.

### v2

@@ Abstract
In high-stakes domains such as healthcare and hiring, the role of machine learning (ML) in decision-making raises significant fairness concerns. This work focuses on Counterfactual Fairness (CF), which posits that an ML model’s outcome on any individual should remain unchanged if they had belonged to a different demographic group. Previous works have proposed methods that guarantee CF. Notwithstanding, their effects on the model’s predictive performance remain largely unclear. To fill this gap, we provide a theoretical study on the inherent trade-off between CF and predictive performance in a model-agnostic manner. We first propose a simple but effective method to cast an optimal but potentially unfair predictor into a fair one with minimal performance degradation. By analyzing the excess risk incurred by perfect CF, we quantify this inherent trade-off. Further analysis on our method’s performance with access to only incomplete causal knowledge is also conducted. Built upon this, we propose a practical algorithm that can be applied in such scenarios. Experiments on both synthetic and semi-synthetic datasets demonstrate the validity of our analysis and methods.
@@ 1 Introduction
Machine learning (ML) has been widely used in high-stakes domains such as healthcare (Daneshjou et al., 2021), hiring (Hoffman et al., 2018), criminal justice (Brennan et al., 2009), and loan assessment (Khandani et al., 2010), bringing with it critical ethical and social considerations. A prominent example is the bias observed in the COMPAS tool against African Americans in recidivism predictions (Brackey, 2019). This issue is particularly alarming in an era where large-scale deep learning models, commonly trained on noisy data from the internet, are increasingly prevalent. Such models, due to their extensive reach and impact, amplify the potential for widespread and systemic biases. This increasing awareness underscores the need for ML practitioners to integrate fairness considerations into their work, extending their focus beyond merely maximizing prediction accuracy (Bolukbasi et al., 2016; Calders and Verwer, 2010; Dwork et al., 2012; Grgic-Hlaca et al., 2016; Hardt et al., 2016). Various fairness notions have been developed, ranging from group-level measures such as group parity (Hardt et al., 2016) to individual-level metrics (Dwork et al., 2012). Recently, there has been a growing interest in approaches based on causal inference, particularly in understanding the causal effects of sensitive attributes such as gender and age on decision-making (Chiappa, 2019; Galhotra et al., 2022; Khademi et al., 2019). This has led to the proposal of Counterfactual Fairness (CF), which states that prediction for an individual in hypothetical scenarios where their sensitive attributes differ should remain unchanged (Kusner et al., 2017). As an individual-level notion agnostic to the choice of similarity measure (Kusner et al., 2017; Rosenblatt and Witter, 2023), CF has recently gained traction (Anthis and Veitch, 2024; Nilforoshan et al., 2022; Makhlouf et al., 2022; Rosenblatt and Witter, 2023).
@@
To theoretically understand the tradeoff between CF and ML performance, we consider a class of invertible causal models and prove that the optimal solution under perfect Counterfactual Fairness (CF) has a simple form w.r.t. to the Bayes optimal classifier and explicitly quantify the excess risk of imposing a perfect CF constraint as has been done for non-causal fairness notions(Zhao and Gordon, 2022; Xian et al., 2023). The optimal predictor under the fairness constraint can be achieved by combining factual and counterfactual predictions using a (potentially unfair) optimal predictor. Next, we quantify the excess risk between the optimal predictor with and without CF constraints. This quantity sheds light on the best possible model, in terms of predictive performance, under the stringent notion of perfect CF. To consider scenarios with incomplete causal knowledge (e.g. unknown causal graph or model), we further study the CF and predictive performance degradation caused by imperfect counterfactual estimations. Inspired by our theoretical findings, we propose a plugin method that leverages a (potentially unfair) pretrained model to achieve a better tradeoff of fairness and predictive performance than the prior methods. Furthermore, we propose a method to improve the pretrained model that accounts for counterfactual estimation errors and can achieve good empirical performance even with limited causal knowledge. We summarize our contributions as follows:
@@
    We propose a CF method that is provably optimal in terms of predictive performance under perfect CF.
@@
    To the best of our knowledge, we are the first to characterize the inherent trade-off of CF and ML performance, which applies to all CF methods.
  3. 3.
@@
  4. 4.
    We empirically demonstrate that our proposed CF methods outperform existing methods in both full and incomplete causal knowledge settings  [^1] Code can be found in https://github.com/inouye-lab/pcf.
@@ 2 Preliminaries
Notation
We use capital letters to represent random variables and lowercase letters to represent the realizations of random variables. Now we define a few variables that will be considered in this work. $A$ represents the sensitive attribute of an individual (e.g., gender), $Y$ represents the target variable to predict, $X$ represents observed features other than $A$ and $Y$, and $U$ represents unobserved confounding variables which are not caused by any observed variables while $a,y,x,u$ represent their realization respectively.
Counterfactual
In this work, we use the framework of Structural Causal Models (SCMs) (Pearl, 2009). A SCM is a triplet $\mathcal{M}=(\bf{U},\bf{V},\mathcal{F})$ where $\bf{U}$ represents exogenous variables (factors outside the model), $\bf{V}$ represents endogenous variables, and $\mathcal{F}$ contains a set of functions $F_{i}$ that map from $U_{i}$ and $Parent(V_{i})$ to $V_{i}$. A counterfactual query asks a question like: what would the value of $Y$ be if $A$ had taken a different value given certain observations? For example, given that a person is a woman and given everything we observe about her performance in an interview, what is the probability of her getting the job if she had been a man? More formally, given a SCM, a counterfactual query can be written as $P(Y_{A=a}|W=w)$. Here $W=w$ is the evidence and $A=a$ in the subscript represents the intervention on $A$. For the general procedure to estimate counterfactuals, please refer to Pearl (2009).
Counterfactual Fairness
Built upon the framework above, we focus on Counterfactual Fairness (CF), which requires the predictors to be fair among factual and counterfactual samples. More formally, it is defined as below
@@ 3.1 Problem Setup
We assume that all data we have is generated by a causal model (Pearl, 2009), and we consider the representative causal graph shown in Figure 1 that has been widely adopted in the fairness literature (Grari et al., 2023; Kusner et al., 2017; Zuo et al., 2023). Our analysis is presented based on binary $A\in\{0,1\}$ given its pivotal importance in the literature (Pessach and Shmueli, 2023) and for the sake of presentation clearness, but our analysis and our method can be naturally extended to multi-class $A$. We first state the main assumptions needed in this section.
$A$$U$$Y$X
Figure 1: Causal graph. $A$ represents sensitive attribute, $Y$ represents the target variable, $U$ represents latent confounders, $X$ represents observed features. Note that the validity of our theoretical analysis holds for all causal models that satisfy the condition given by Assumption 3.1. It is not restricted to this specific graph.
@@
The first assumption is very common in the fairness literature. While the invertibility assumption might be restrictive in certain scenarios, it simplifies the theoretical analysis and has been adopted in recent works on counterfactual estimation (Nasr-Esfahany et al., 2023; Zhou et al., 2024). We expect that exact invertibility is not required in practice but rather only strong mutual information between $X$ and $U$ given $A$ would be sufficient. Further, we empirically validate the effectiveness of our method after relaxing invertibility in the experiment section. To facilitate our discussion, we first define $F_{X}$ as the mapping between $X$ and $(U,A)$, i.e., $X=F_{X}(U,A)$. According to our second assumption, $F_{X}(\cdot,a)$ is an invertible function, i.e., $\exists{F_{X}^{*}}^{-1},{F_{X}^{*}}^{-1}(x,a)={F_{X}^{*}}^{-1}(F_{X}^{*}(u,a),a)=u,\forall(x,a)$. This assumption simplifies the counterfactual estimation of $X$ for different values of $A$ into a deterministic function. In our context, the counterfactual query is specifically $p(X_{a^{\prime}}|X=x,A=a)$, which simplifies to a Dirac delta at a single value given the invertibility assumption. Thus, we introduce the concept of a deterministic counterfactual generating mechanism (CGM), denoted as $x_{a^{\prime}}=G(x,a,a^{\prime})$. Also, in our case, we will assume $A$ is binary so that $a$ and it’s counterfactual $a^{\prime}$ can be written as $1-a$. All proofs can be found in Appendix A. Given this setup, the following lemma characterizes the perfect CF constraint on $\phi$.
@@
The proof is straightforward from the definition of TE. Notably, this lemma helps disambiguate the question of whether counterfactual fairness is a distribution- or individual-level requirements as raised in Plecko and Bareinboim (2022). In our setup, they are equivalent due to invertibility between $X$ and $U$ given $A$.
@@ 3.2 Optimal Counterfactual Fairness and Inherent Trade-off
Given the complete knowledge of the causal model, it is viable to satisfy the perfect CF constraint (Kusner et al., 2017; Zuo et al., 2023). However, these methods are known to result in empirical degradation of ML models’ performance, raising critical concerns about the fairness-utility trade-offs. Moreover, it is still unknown to what extent the ML model performance has to be affected in order to achieve perfect CF. In this section we provide a formal study on this to close the gap. Our solution consists of two steps. First, we propose a simple yet effective method that is provably optimal under the constraint of perfect CF. Next, we characterize the inherent trade-off between CF and predictive performance by checking the excess risk compared to a Bayes optimal (unfair) predictor. Our result shows that the inherent trade-off is dominated by the dependency between $Y$ and $A$, echoing previous analysis on non-causal based fairness notions (Chzhen et al., 2020; Xian et al., 2023). For brevity, we refer to a Bayes optimal predictor as “optimal”, and a model that satisfies perfect CF as “fair”.
@@
This result suggests that, if we have to access to ground truth counterfactuals, a simple algorithm using a (potentially) unfair model could achieve strong fairness and accuracy. Built upon the above result, we are ready to characterize the inherent trade-off between CF and model performance by the following theorem.
@@
Remarkably, the excess risks are completely characterized by the inherent dependency between $Y$ and $A$ as determined by the underlying causal mechanism, similar to non-causal based group fairness (Chzhen et al., 2020; Xian et al., 2023). Moreover, they lower bound the excess risk of all possible predictors in order to achieve perfect CF.
@@ 3.3 Method with Incomplete Causal Knowledge
In this section, we aim to address CF in the scenario where causal knowledge is limited. Inspired by Theorem 3.3, we first present a simple plugin method as summarized in Algorithm 1. For regression tasks, $\hat{\mu}$ is the final output, and for classification tasks, $\hat{\mu}$ represents the probability of $Y=1$, i.e., $p(Y=1|X=x,A=a)=\mathbb{E}[Y|X=x,A=a]$. It is noteworthy that PCF is agnostic to the training of predictor $\phi$ that can be determined by the user freely. In fact, with access to the oracle CGM $G^{*}$, then PCF would achieve perfect CF as proved in the next result.
@@
Note that this proposition only requires access to ground truth $G^{*}$ and holds valid for any pretrained predictor $\phi$. If $\phi$ is further accurate, then the corresponding PCF is able to achieve high accuracy as well, which is empirically validated in the experiments.
@@ 3.3.1 Given estimated $G$
Acquiring counterfactuals in practice can be a challenging task and could lead to estimation errors. In this section we provide a theoretical analysis on this. Specifically, the theorem below bounds the TE and excess risk due to the use of estimated counterfactuals.
@@
Given an optimal predictor $\phi^{*}(x,a)$, suppose it is L-lipschitz continuous in $x$, and the counterfactual estimation error is bounded, i.e.,
@@
for some $\varepsilon\geq 0$, where $G^{*}$ and $\hat{G}$ represent the ground truth and estimated CGMs respectively. Then, the total effect (TE) of Algorithm 1 based on $\hat{G}$ is bounded by $L\varepsilon$. Moreover, for squared $L_{2}$ loss, the excess risk is bounded by $\sigma_{A}^{2}L^{2}\varepsilon^{2}+2\sigma_{A}^{2}L\varepsilon\mathbb{E}_{U}[|\mathbb{E}[Y\mid U=u,A=1]-\mathbb{E}[Y\mid U=u,A=0]|]$, and for cross-entropy loss  [^3] Here we assume the logits are L-lipschitz continuous in $x$., the excess risk is bounded by $L\varepsilon$.
Note that $\sigma_{A}$ and $\mathbb{E}_{U}[|\mathbb{E}[Y\mid U=u,A=1]-\mathbb{E}[Y\mid U=u,A=0]|]$ are inherent characteristic of the underlying mechanism and is independent of the counterfactual estimation. This suggests that if the counterfactuals are not too far away and $\phi^{*}$ is smooth, then fairness and prediction performance will not be significantly affected. In practice, CGM in Algorithm 1 can be obtained using counterfactual estimation methods, as discussed in Section 6.
@@ 3.3.2 Given estimated $G$ and $\phi$
The proof is straightforward since the original proof in Theorem 3.6 does not use any characteristic of optimality. To achieve good predictive performance, a natural approach is to train $\phi$ on the observed data via Empirical Risk Minimization (ERM), which should fit the predictor well given sufficient samples and a reasonable predictor class. However, ERM can only approximate Bayes optimality within the support of the training data. Outside this support, its performance can deteriorate significantly, as extensively studied in areas such as Domain Adaptation (Farahani et al., 2021) and Domain Generalization (Zhou et al., 2022a). Consequently, when integrated with an approximate $G$, we may encounter the issue where $p(\hat{X}_{A=1-a}|A=a)\neq p(X|A=1-a)$, inducing a distribution shift problem (Note that these would be equal given the graph in Figure 1 and Pearl’s rules) (Kulinski and Inouye, 2023). To mitigate this, we suggest improving $\phi$ on the estimated counterfactual distribution. More formally we define the following objective called Counterfactual Risk Minimization (CRM):
@@
This can be achieved either by augmenting the original training dataset or by fine-tuning with estimated counterfactual samples. The choice between training from scratch or fine-tuning depends on the scale of the experiment and computational constraints. It is important to note that the $Y$ corresponding to the estimated counterfactual should remain the same with that of the factual samples. While the optimal prediction for the counterfactual may differ from that of the original data, under the constraint of perfect CF, a predictor is required to predict the same outcome to counterfactual pairs. Hence, the optimal solution will change. This is exactly what causes the excess risk we characterized in Theorem 3.4. Furthermore, we can prove that, given the ground truth $G$, CRM yields the same optimal solution as PCF. Since this result is dependent on the ground truth $G$, we provide a more formal statement in Appendix B to ensure consistency.